# DFT-s-OFDM-based On-Off Keying for Low-Power Wake-Up Signal

### v1

@@ VII Conclusions
In [14], the subcarrier coefficients are obtained as [14, Eq. (8)] $\tilde{{\bf x}}_{LS}=(\tilde{{\bf F}}^{H}\tilde{{\bf F}})^{-1}\tilde{{\bf F}}^{H}{\bf b}_{m}$ where ${\bf b}_{m}\in\{0,1\}^{N_{\rm fft}\times 1}$ contains an upsampled version of the bits string $b[m]$ matching the OFDM IFFT size (i.e. $b_{s}^{\rm LS}[n]$ in (23)), $\tilde{{\bf F}}=\frac{1}{N_{\rm fft}}[e^{j\frac{2\pi mn}{N_{\rm fft}}}]_{\begin{subarray}{c}n\in\mathcal{K}_{c}\\ 0\leq m<N_{\rm fft}\end{subarray}}$ is a truncated $N_{\rm fft}\times N_{\rm sc}$ IDFT matrix whose columns have been restricted to the shifted indexes $\mathcal{K}_{c}=(\mathcal{K}-k_{c})\;\text{mod}\,N_{\rm fft}$ with $\mathcal{K}=f_{0}+\{0,\ldots,N_{\rm sc}-1\}$ corresponding to the LP-WUS subcarrier allocation and $k_{c}$ being the center index of the subset $\mathcal{K}$. The value of the middle subcarrier $k_{c}$ is not explicitly given in [14], so we will conventionally take $k_{c}=f_{0}+\lfloor\frac{N_{\rm sc}}{2}\rfloor$ but the alternative $k_{c}=f_{0}+\lceil\frac{N_{\rm sc}}{2}\rceil$ could also be selected.
In 3GPP study on LP-WUS [2], this construction was discussed as a specific type of precoding, different from DFT, but it is actually simply a DFT where only some outputs are retained. Indeed, the regularization matrix $(\tilde{{\bf F}}^{H}\tilde{{\bf F}})^{-1}$ is given following the generic least-squares formulation, however here by construction this simplifies to $(\tilde{{\bf F}}^{H}\tilde{{\bf F}})^{-1}=N_{\rm fft}{\bf I}_{N_{\rm sc}}$ such that this method reduces to $\tilde{{\bf x}}_{LS}=\bar{{\bf F}}{\bf b}_{m}$ where $\bar{{\bf F}}=[e^{-{\rm j}\frac{2\pi mn}{N_{\rm fft}}}]_{\begin{subarray}{c}0\leq n<N_{\rm fft}\\ m\in\mathcal{K}_{c}\end{subarray}}$ is a truncated $N_{\rm sc}\times N_{\rm fft}$ DFT matrix where only $N_{\rm sc}$ rows/outputs corresponding of subcarriers indices in $\mathcal{K}_{c}$ are preserved. Note that the selected DFT output indices $\mathcal{K}_{c}$ corresponds simply to the baseband subcarrier indices $\mathcal{K}_{c}=\{-\lfloor\frac{N_{\rm sc}}{2}\rfloor,\ldots,\lfloor\frac{N_{\rm sc}-1}{2}\rfloor\}\;\text{mod}\,N_{\rm fft}$. We encountered in engineer circle that this indexing notation in [14] may be misunderstood by some readers. This is reflected in [6] where the description of [14] is using instead the truncated indices $\{0,\ldots,N_{\rm sc}-1\}$ which does not correspond to [14] and so the generated LS waveform therein has much less a rectangular shape than the waveform obtained from [14] should have.
@@
From (15), the average power of the subcarrier coefficients is direclty given by $E[|D_{0}[k]|^{2}]=|R_{0}[k]|^{2}E[|B[k]|^{2}]$.
@@
$E[|\mathcal{B}[k]|^{2}]$ $=$ $\frac{N_{\rm bit}}{2}+\frac{1}{4}\sum_{\begin{subarray}{c}n,m=0\\ n\neq m\end{subarray}}^{N_{\rm bit}-1}1=\frac{N_{\rm bit}}{4}+\frac{1}{4}\sum_{n,m=0}^{N_{\rm bit}-1}1$ (43)
@@
Puttiing all together, we have
$E[|B[k]|^{2}]=\begin{cases}\frac{N_{\rm bit}(N_{\rm bit}+1)}{4}\quad\quad{\rm for}\;\;k=(0\,{\rm mod}\,N_{\rm bit})\\ \frac{N_{\rm bit}}{4}\quad{\rm otherwise}\end{cases}.$ (45)
-E2 Half of the Total Power on a Regular Comb
For a $N$-length sequence $s=(s[0],\ldots,s[N-1])$, let us write its total power as $\|s\|^{2}=\sum_{n=0}^{N-1}|s[n]|^{2}$. The $N$-DFT is a unitary operation if normalized by $1/\sqrt{N}$, so for the DFT sequence $S[k]=\sum_{n=0}^{N-1}s[n]e^{-{\rm j}2\pi\frac{nk}{N}}$, its total power is $\|S\|^{2}=N\|s\|^{2}$.
With Manchester coding, the power of the bits string is always $\|b\|^{2}=N_{\rm bo}$, so the total power of the modulating symbols (5) is $\|d\|^{2}=N_{\rm bo}\|r_{0}\|^{2}$, and the total power of subcarrier coefficients (6) is $\|D\|^{2}=N_{\rm bo}N_{\rm pulse}\|r_{0}\|^{2}$. Then, the share of power of these subcarrier coefficients on the comb $\{k=k^{\prime}N_{\rm bit}\}_{k^{\prime}=0}^{N_{\rm seg}-1}$ is
$\sum_{k^{\prime}=0}^{N_{\rm seg}-1}|D_{0}[k^{\prime}N_{\rm bit}]|^{2}\!\!\!\!$ $=$ $\!\!\!\!\!\sum_{k^{\prime}=0}^{N_{\rm seg}-1}|R_{0}[k^{\prime}N_{\rm bit}]|^{2}|B[k^{\prime}N_{\rm bit}]|^{2}$ (46)
@@
where the second equality follows from (28) and $|B[k^{\prime}N_{\rm bit}]|^{2}=|\mathcal{B}[0]|^{2}=N_{\rm bo}^{2}$ due to the repetitive structure of $B[k]$. Since $\frac{\sum_{k=0}^{N_{\rm seg}-1}|D[kN_{\rm bit}]|^{2}}{\|D\|^{2}}=\frac{N_{\rm bo}N_{\rm seg}}{N_{\rm pulse}}=\frac{1}{2}$, half of the total power of the subcarrier coefficients is allocated on this comb.
Without Manchester coding, the average total power of the bits string is $E[\|b\|^{2}]=N_{\rm bit}/2$, so on average we have also $E[\|d\|^{2}]=N_{\rm bit}/2\|r_{0}\|^{2}$, and $E[\|D\|^{2}]=\|r_{0}\|^{2}N_{\rm pulse}N_{\rm bit}/2$. The share of average power on the comb is then
$\sum_{k^{\prime}=0}^{N_{\rm seg}-1}E[|D_{0}[k^{\prime}N_{\rm bit}]|^{2}]$ $=$ $\frac{N_{\rm bit}(N_{\rm bit}+1)}{4}N_{\rm seg}\|r_{0}\|^{2}$
and the ratio is $\frac{\sum_{k=0}^{N_{\rm seg}-1}E[|D[kN_{\rm bit}]|^{2}]}{E[\|D\|^{2}]}=\frac{(N_{\rm bit}+1)}{2N_{\rm bit}}>\frac{1}{2}$.
@@
By definition $R_{0}[k]=\sum_{l=0}^{N_{\rm seg}-1}r_{0}[l]e^{-{\rm j}\frac{2\pi kl}{N_{\rm pulse}}}$. Now assume that the DFT of $r_{0}[l]$ is $R^{\prime}_{0}[h]$ such that $r_{0}[l]=\frac{1}{N_{\rm seg}}\sum_{h=0}^{N_{\rm seg}-1}R^{\prime}_{0}[h]e^{{\rm j}\frac{2\pi lh}{N_{\rm seg}}}$. Combining them and using the fact that $N_{\rm pulse}=N_{\rm seg}N_{\rm bit}$, we get
@@
If $k=xN_{\rm bit}$ with $x$ integer,
$R_{0}[xN_{\rm bit}]=\frac{1}{N_{\rm seg}}\sum_{h=0}^{N_{\rm seg}-1}R^{\prime}_{0}[h]\sum_{l=0}^{N_{\rm seg}-1}e^{-{\rm j}\frac{2\pi l(x-h)}{N_{\rm seg}}}.$ (48)
The inner sum above is equal to zero if $h\neq x$ and otherwise to $N_{\rm seg}$, so we have $R_{0}[xN_{\rm bit}]=R^{\prime}_{0}[x]$ which gives the first line of (28).
If $k\neq xN_{\rm bit}$, using the sum exponential formula,
@@
and so we get
$R_{0}[k]=\frac{1}{N_{\rm seg}}\sum_{h=0}^{N_{\rm seg}-1}R^{\prime}_{0}[h]\frac{\sin\left(\pi\frac{(k-N_{\rm bit}h)}{N_{\rm bit}}\right)}{\sin\left(\pi\frac{(k-N_{\rm bit}h)}{N_{\rm pulse}}\right)}e^{-{\rm j}\frac{\pi(N_{\rm seg}-1)}{N_{\rm pulse}}(k-N_{\rm bit}h)}$ (50)
which leads to the second line of (28).

### v2

@@ Acknowledgments
In [14], the subcarrier coefficients are obtained as [14, Eq. (8)] $\tilde{{\bf x}}_{LS}=(\tilde{{\bf F}}^{H}\tilde{{\bf F}})^{-1}\tilde{{\bf F}}^{H}{\bf b}_{m}$ where ${\bf b}_{m}\in\{0,1\}^{N_{\rm fft}\times 1}$ contains an upsampled version of the bits string $b[m]$ matching the OFDM IFFT size (i.e. $b_{\rm r}^{\rm LS}[n]$ in (25)), $\tilde{{\bf F}}=\frac{1}{N_{\rm fft}}[e^{j\frac{2\pi mn}{N_{\rm fft}}}]_{\begin{subarray}{c}n\in\mathcal{K}_{c}\\ 0\leq m<N_{\rm fft}\end{subarray}}$ is a truncated $N_{\rm fft}\times N_{\rm sc}$ IDFT matrix whose columns have been restricted to the shifted indexes $\mathcal{K}_{c}=(\mathcal{K}-k_{c})\;\text{mod}\,N_{\rm fft}$ with $\mathcal{K}=f_{0}+\{0,\ldots,N_{\rm sc}-1\}$ corresponding to the LP-WUS subcarrier allocation and $k_{c}$ being the center index of the subset $\mathcal{K}$. The value of the middle subcarrier $k_{c}$ is not explicitly given in [14], so we will conventionally take $k_{c}=f_{0}+\lfloor\frac{N_{\rm sc}}{2}\rfloor$ but the alternative $k_{c}=f_{0}+\lceil\frac{N_{\rm sc}}{2}\rceil$ could also be selected.
In 3GPP LP-WUS study [2], this construction was often discussed as a specific type of precoding, different from DFT, while it is actually simply a large DFT where only some outputs are retained. Indeed, the regularization matrix $(\tilde{{\bf F}}^{H}\tilde{{\bf F}})^{-1}$ is given following the generic least-squares formulation, however here this simplifies to $(\tilde{{\bf F}}^{H}\tilde{{\bf F}})^{-1}=N_{\rm fft}{\bf I}_{N_{\rm sc}}$ such that this method reduces to $\tilde{{\bf x}}_{LS}=\bar{{\bf F}}{\bf b}_{m}$ where $\bar{{\bf F}}=[e^{-{\rm j}\frac{2\pi mn}{N_{\rm fft}}}]_{\begin{subarray}{c}0\leq n<N_{\rm fft}\\ m\in\mathcal{K}_{c}\end{subarray}}$ is a truncated $N_{\rm sc}\times N_{\rm fft}$ DFT matrix where only $N_{\rm sc}$ rows/outputs corresponding of subcarriers indices in $\mathcal{K}_{c}$ are preserved. Note that the selected DFT output indices $\mathcal{K}_{c}$ corresponds simply to the baseband subcarrier indices $\mathcal{K}_{c}=\{-\lfloor\frac{N_{\rm sc}}{2}\rfloor,\ldots,\lfloor\frac{N_{\rm sc}-1}{2}\rfloor\}\;\text{mod}\,N_{\rm fft}$. We encountered that this indexing notation in [14] can be misunderstood by some readers. This is reflected in [6] which is built on improving the LS approach from [14] but their description of it corresponds to use $\mathcal{K}_{c}=\{0,\ldots,N_{\rm sc}-1\}$ instead. Hence, the “LS approach” in [6] has a much poorer rectangular shape than it should have.
@@
We start with the average power of the subcarrier coefficients given by $E[|D_{0}[k]|^{2}]=|R_{0}[k]|^{2}E[|B[k]|^{2}]$ from (16).
@@
$E[|\mathcal{B}[0]|^{2}]$ $=$ $\frac{N_{\rm bit}}{2}+\frac{1}{4}\sum_{\begin{subarray}{c}n,m=0\\ n\neq m\end{subarray}}^{N_{\rm bit}-1}1=\frac{N_{\rm bit}}{4}+\frac{1}{4}\sum_{n,m=0}^{N_{\rm bit}-1}1$ (47)
@@
Putting all together, we have
$E[|B[k]|^{2}]=\begin{cases}\frac{N_{\rm bit}(N_{\rm bit}+1)}{4}\quad\quad{\rm for}\;\;(k\,{\rm mod}\,N_{\rm bit})=0\\ \frac{N_{\rm bit}}{4}\quad{\rm otherwise}\end{cases}.$ (49)
-E2 Half Power Allocation on a Regular Comb
For a $N$-length sequence $s=(s[0],\ldots,s[N-1])$, let us write its power as $\|s\|^{2}=\sum_{n=0}^{N-1}|s[n]|^{2}$. The $N$-DFT is a unitary operation if normalized by $1/\sqrt{N}$, so the DFT sequence $S[k]=\sum_{n=0}^{N-1}s[n]e^{-{\rm j}2\pi\frac{nk}{N}}$ has power $\|S\|^{2}=N\|s\|^{2}$.
With Manchester coding, the bits string always satisfies $\|b\|^{2}=N_{\rm bo}$, so the modulating symbols (5) fulfills $\|d\|^{2}=N_{\rm bo}\|r_{0}\|^{2}$, and the sum power of subcarrier coefficients (6) is $\|D\|^{2}=N_{\rm bo}N_{\rm pulse}\|r_{0}\|^{2}$. The share of this sum power on the comb $\mathcal{C}=\{k^{\prime}N_{\rm bit}\}_{k^{\prime}=0}^{N_{\rm seg}-1}$ can be computed to be
$\sum_{k\in\mathcal{C}}|D_{0}[k]|^{2}\!\!\!\!$ $=$ $\!\!\!\!\!\sum_{k^{\prime}=0}^{N_{\rm seg}-1}|R_{0}[k^{\prime}N_{\rm bit}]|^{2}|B[k^{\prime}N_{\rm bit}]|^{2}$ (50)
@@
where the second equality follows from (30) and $|B[k^{\prime}N_{\rm bit}]|^{2}=|\mathcal{B}[0]|^{2}=N_{\rm bo}^{2}$ due to the repetitive structure of $B[k]$. Since $\frac{\sum_{k\in\mathcal{C}}|D[k]|^{2}}{\|D\|^{2}}=\frac{N_{\rm bo}N_{\rm seg}}{N_{\rm pulse}}=\frac{1}{2}$, this comb always carries half of the sum power.
Without Manchester coding, the average power of the bits string is $E[\|b\|^{2}]=N_{\rm bit}/2$, so on average we also have $E[\|d\|^{2}]=N_{\rm bit}/2\|r_{0}\|^{2}$, and $E[\|D\|^{2}]=\|r_{0}\|^{2}N_{\rm pulse}N_{\rm bit}/2$. From (49), the share of average power on the comb becomes
$\sum_{k\in\mathcal{C}}E[|D_{0}[k]|^{2}]=\frac{N_{\rm bit}(N_{\rm bit}+1)}{4}N_{\rm seg}\|r_{0}\|^{2}$ (51)
and the ratio $\frac{\sum_{k\in\mathcal{C}}E[|D[k]|^{2}]}{E[\|D\|^{2}]}=\frac{(N_{\rm bit}+1)}{2N_{\rm bit}}>\frac{1}{2}$.
@@
We have $R_{0}[k]=\sum_{l=0}^{N_{\rm seg}-1}r_{0}[l]e^{-{\rm j}\frac{2\pi kl}{N_{\rm pulse}}}$, while by definition $r_{0}[l]=\frac{1}{N_{\rm seg}}\sum_{h=0}^{N_{\rm seg}-1}R^{\prime}_{0}[h]e^{{\rm j}\frac{2\pi lh}{N_{\rm seg}}}$. Combining them and using the fact that $N_{\rm pulse}=N_{\rm seg}N_{\rm bit}$, we get
@@
If $k=k^{\prime}N_{\rm bit}$ with $k^{\prime}$ integer,
$R_{0}[k^{\prime}N_{\rm bit}]=\frac{1}{N_{\rm seg}}\sum_{h=0}^{N_{\rm seg}-1}R^{\prime}_{0}[h]\sum_{l=0}^{N_{\rm seg}-1}e^{-{\rm j}\frac{2\pi l(k^{\prime}-h)}{N_{\rm seg}}}.$ (53)
The inner sum above is equal to zero if $h\neq k^{\prime}$ and otherwise to $N_{\rm seg}$, so we have $R_{0}[k^{\prime}N_{\rm bit}]=R^{\prime}_{0}[k^{\prime}]$ which gives the first line of (30).
If $k\neq k^{\prime}N_{\rm bit}$, using the sum exponential formula,
@@
and so we get the second line of (30).